Circular No. 5912

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993ai IN UGC 3483 Jean Mueller reports her discovery of a supernova (blue mag ∼ 18)

located 23′′.5 west of the nucleus of UGC 3483 (α = 6^{h}29^{m}.3, δ = +51^{o}23′, equinox 1950.0). The object was found on a plate taken on Dec. 10 UT by D. J. Griffith, J. D. Mendenhall, and Mueller with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey.

A. V. Filippenko and L. C. Ho, University of California at Berkeley; and W. L. W. Sargent, California Institute of Technology, report that an uncalibrated spectrum (range 350-750 nm) obtained on Dec. 21 with the 5-m Hale reflector at Palomar Observatory suggests that the supernova is of type Ia, roughly 1 month past maximum brightness. A redshift of ∼ 0.034 was measured for the parent galaxy from the Na I D absorption line.

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence,

write: “CCD photometry of this repeating transient (IAUC 5842, 5901) obtained with the 1.2-m Haute-Provence telescope on Dec. 17.0 UT gives V = 15.38, B–V = +0.37, V–I_{c} = +0.67. Observations on Dec. 18 give essentially the same colors at a level 0.05 mag fainter in V. Although the source is now at a level comparable with that reached during its brief August outburst, it is slightly bluer than when measured on Aug. 16 with the same instrument (V = 15.36, B–V = +0.42, V–I_{c} = +0.71). During that brief outburst, which started on Aug. 10, maximum brightness was reached on Aug. 14 (V = 15.2), and the phase of rapid decline started around Aug. 30, with the object reaching V = 16.8 on Sept. 6.”

NOVA CASSIOPEIAE 1993 Photoelectric photometry: Dec. 17.77 UT, V = 5.76 ± 0.03, B–V =

+0.69 ± 0.04 (H. Mikuž, Ljubljana, Slovenia; cf. IAUC 5905); 17.86, V = 5.69 (E. Poretti, Merate Observatory; comparison star HR 8985, V = 6.24).

Further visual magnitude estimates (cf. IAUC 5910): Dec. 18.23 UT, 5.8 (E. Waagen, Cambridge, MA); 18.73, 6.1 (S. Meister, Bulach, Switzerland); 19.07, 6.7 (P. Collins, Scottsdale, AZ); 19.666, 7.3 (L. Szentaskó, Budapest, Hungary); 19.73, 6.9 (G. Comello, Groningen, The Netherlands); 20.24, 6.4 (R. Royer, Lakewood, CA).

1993 December 21 Daniel W. E. Green